Suppose that a salesperson succeeds in making a sale with three in every 4 potential customers contacted. a particular week 5 new potential customers are conducted.

What is the probability read the salesperson will succeed in making at least 4 sales?


1
Expert's answer
2021-05-24T03:46:48-0400

"P=\dfrac{3}{4}=0.75"


Given distribution is a binomial distribution

"X\sim bin(5,0.75)"


 Probability that the salesperson will succeed in making at least 4 sales-


"P(X\ge 4)=1-P(X>4)"


"\\=1-P(X=5)\\[9pt]=1-^5C_5(0.75)^5(0.25)^0\\[9pt]=1-0.2373=0.7626"